A firm has a monopoly in the market for match boxes. In order to produce Q match boxes, it cost a firm C(Q)=2Q^2
A) Find the marginal cost of producing a matchbox for this firm?
1
Expert's answer
2020-11-27T12:46:38-0500

"\bold {Answer}"

"MC = 4Q"


"\bold {Solution}"

Marginal cost is found by differential calculus, given the total cost function.

"MC = \dfrac {d}{dQ} (TC)"


"= \dfrac {d}{dQ} [C(Q)]"


"= \dfrac {d}{dQ} (2Q^2)"


"= 4Q"


Therefore, the marginal cost is "\$4Q"